OpenStudy (anonymous):

Express the set using set buider notation. Use inequality notation to express the consition x must meet in order to be a member of the set. A={200,201,202,....2000

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (zzr0ck3r):

\[\{x\in\mathbb{Z} \ | \ a\le x\le b \}\ \]

OpenStudy (zzr0ck3r):

what do you think a and b are?